OpenStudy (anonymous):

Sarah’s dad tells her she can have the greater of the following two amounts for her allowance (in dollars): 2 times the difference of 5 and 1 5 times 1 minus 2 What is Sarah’s allowance, in dollars? Numerical Answers Expected!

OpenStudy (anonymous):

2 times the difference of 5 and 1 = 2(5-1)=2(4)=8 5 times 1 minus 2 = 5(1) -2 = 5-2 =3 So, $8